### v1

# DepsRAG: Towards Managing Software Dependencies using Large Language Models

Abstract.
Managing software dependencies is a crucial maintenance task in software development and is becoming a rapidly growing research field, especially in light of the significant increase in software supply chain attacks. Specialized expertise and substantial developer effort are required to fully comprehend dependencies and reveal hidden properties about the dependencies (e.g., number of dependencies, dependency chains, depth of dependencies) .
Recent advancements in Large Language Models (LLMs) allow the retrieval of information from various data sources for response generation, thus providing a new opportunity for uniquely managing software dependencies. To highlight the potential of this technology, we present DepsRAG, a proof-of-concept Retrieval Augmented Generation (RAG) approach that constructs direct and transitive dependencies of software packages as a Knowledge Graph (KG) in four popular software ecosystems. DepsRAG can answer user questions about software dependencies by automatically generating necessary queries to retrieve information from the KG, and then augmenting the input of LLMs with the retrieved information. DepsRAG can also perform Web search to answer questions that the LLM cannot directly answer via the KG. We identify tangible benefits that DepsRAG can offer and discuss its limitations. [^1] DepsRAG demo and implementation are available here: https://shorturl.at/npyIZ.
copyright: none
1. Introduction
Software development increasingly relies on code reuse, which can be accomplished through software packages. Utilizing third-party packages to build software can significantly improve quality and productivity (Lim, 1994; Mohagheghi et al., 2004). These packages, dependencies, and their maintainers have formed large software ecosystems (Sonatype, 2021). Managing dependencies between packages is an emerging and important problem (Artho et al., 2012; Bogart et al., 2016; Decan et al., 2019), especially as open-source and third-party components have become widely adopted in software development. This new trend highlights the importance of software supply chain security, which is becoming more critical over time.
Today, developers are expected to manage their software dependencies (Kula et al., 2018). However, in a recent survey, developers reported that they are not confident in their current dependency management practices (Tidelift, [n. d.]). While developers often have full control over which dependency packages and versions to use, they have no control over the dependencies of the used packages. Such transitive dependencies, which often amount to more than direct dependencies, may have various vulnerabilities that put software projects at risk. To mitigate these risks, we developed DepsRAG to support practitioners and software developers understand the dependencies in their software supply chain. DepsRAG facilitates constructing and analyzing software dependencies for four popular software ecosystems. It also answers key questions about the dependency graph, including the number of dependencies, the depth of the dependency graph, and the number of paths/chains. The use of LLM enables DepsRAG to autonomously generate queries for KGs, unlike prior methods demanding hardcoded or user-generated queries (Dann et al., 2023; Litzenberger et al., 2023; Yamaguchi et al., 2014), thereby enhancing scalability. This retrieved information enhances the user’s question, referred to as Retrieval Augmented Generation (RAG), enabling accurate LLM responses. DepsRAG works in two phases. In the first phase, the user provides the name and ecosystem of the project. Then it transitively constructs the dependencies as a KG. In the second phase, the user can chat with DepsRAG to answer questions about the dependencies. DepsRAG is equipped with several tools to facilitate its functionality to access the KG, automatically generate queries, search the Web, and visualize the dependency graph.
2. Background
2.1. Dependency Graphs
Dependency graphs are a fundamental concept in software engineering and are used to represent dependencies among various entities of a software system. These entities can be classes, functions, modules, packages, or even larger components. Each node in the graph represents an entity, and each directed edge between two nodes indicates that one entity depends on another. This means that changes in one entity can affect the other.
Litzenberger et al. (Litzenberger et al., 2023) proposed a framework the defines three levels of granularity for dependency representation: package-to-package, artifact-to-package, and artifact-to-artifact. Düsing and Hermann (Düsing and Hermann, 2022) constructed a dependency graph at the artifact-to-package level and leveraged these graphs to investigate the propagation of vulnerabilities in different software repositories. Benelallam et al. (Benelallam et al., 2019) constructed the Maven Dependency Graph to answer high-level research questions about artifact releases, evolution, and usage trends over time. Dependency graphs have also been used to build tools for program understanding (Falke et al., 2005; Musco et al., 2017).
2.2. Knowledge Graphs (KGs)
Knowledge graphs (KGs) organize information in a structured format, capturing relationships between real-world entities and making them comprehensible to humans and machines (Agrawal et al., 2023). In a KG, data are organized as triplets (head entity, relation, tail entity), where the relation is the relationship between these two entities, such as (”Steven Jobs”, ”owns”, ”Apple”). More formally, KGs store structured knowledge as a collection of triples $\mathcal{KG}=\{(e_{i},r,e_{j})\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}\}$, where $\mathcal{E}$ and $\mathcal{R}$ denote the set of entities and relations, respectively. KGs are created by describing entities and entity relationships, known as graph schema. KGs are useful for a variety of applications, such as question-answering (Hao et al., 2017), information retrieval (Xiong et al., 2017), recommender systems (Zhang et al., 2016), cybersecurity (Agrawal, 2023; Yamaguchi et al., 2014), and natural language processing (Yang and Mitchell, 2017).
2.3. Retrieval-Augmented Generation (RAG)
RAG (Lewis et al., 2020; Guu et al., 2020) augments the input questions for LLMs with information retrieved from documents, knowledge graphs, or vector databases, and is considered a useful method for addressing LLM hallucinations. As such, RAG provides extra knowledge sources for LLMs and greatly improves their performance in a variety of tasks, especially in the knowledge-intensive ones.
RAG applications leverage tools and function calls to perform their job. This feature is introduced by LLM multi-agent frameworks (i.e., LangChain (Chase, 2022) and Langroid (lan, 2024)) and recent versions of OpenAI models. [^2] https://platform.openai.com/docs/guides/text-generation/function-calling This feature enables users to provide function descriptions to the LLM. The LLM responds with a structured response, such as a JSON file containing the required arguments to invoke any available functions. These functions serve as action executors and provide the option to supply APIs that the LLM can query to obtain essential information for responding to users. For example, when a user asks about the current weather in a specific city and provides the LLM with a weather API call description, the LLM can augment its response with information retrieved from the API. This facilitates a seamless and efficient interaction between the user and the LLM and provides rich responses that avoid hallucinations.
3. Approach
Figure 1. DepsRAG high-level architecture
DepsRAG consists of two retrievers whose responses augment the prompt with additional information, as illustrated in Figure 1.
  • KG Retriever. This component retrieves information from the KG. It translates the user’s question in natural language to the query language used by the KG, executes the query, and sends back the response to DepsRAG. This component is equipped with two tools to perform these tasks: GraphSchemaTool and CypherQueryTool. The former tool provides the KG schema to DepsRAG, which is needed to generate the query with the right entity names and relationships. The other tool receives the query generated by the LLM, executes it, and returns the response.
  • Web Search Retriever. This component acts as a failover mechanism and answers when no answer is obtained from the KG. It also provides answers beyond the scope of the KG. For example, the user could ask DepsRAG if one of the packages in the dependency KG is vulnerable. This component leverages WebSearchTool to perform a Web search and retrieve top-3 results, and then the LLM can extract the most relevant results.
Figure 3 outlines the prompt used by DepsRAG. This prompt consists of three parts. R defines the role of DepsRAG. $\textrm{p}_{1}$ and $\textrm{p}_{2}$ describe the instructions for the two phases of DepsRAG:
  • $\textrm{p}_{1}$
    Dependency KG Construction. The user provides the name of the package, the version, and the ecosystem, and then DepsRAG constructs the dependency graph following the graph schema illustrated in Figure 2. According to this schema, the nodes in the KG belong to the entity Package, which has only two properties, the package name and version. For entity relationships, there is only one relation depends_on. DepsRAG leverages ConstructKGTool for constructing the KG. Suppose the user did not provide the version and ecosystem of the package. In this case, DepsRAG triggers the Web Search Retriever to find the version and ecosystem based on the provided package name. Figure 3 illustrates the workflow of this phase to construct the dependency graph of the OpenAI package in the PyPI ecosystem.
  • $\textrm{p}_{2}$
    QA Chatting. After constructing the dependency KG for the provided package name, the user can ask any questions about the KG. In this phase, the KG Retriever and Web Search Retriever are employed to augment user questions with additional information. Figure 3 illustrates the workflow of this phase.
Figure 3: DepsRAG Instructions R Define LLM Role: You are an expert in software dependency graphs and analyzing them using Neo4j.
$\textrm{p}_{1}$ Phase-1 Task Instructions: FIRST, I’ll give you the name of the package that I want to analyze.
THEN, you can use the ”WebSearchTool” to find out information about a package, such as version number and package type (PyPI, NPM, Cargo, or GO). If you cannot get this information, ask me and I can tell you.
DON’T forget to include the package name in your questions. After receiving this information, ensure to get the package version type.
THEN ask the user if they want to construct the dependency graph, and if so, use ”ConstructKGTool” to construct the dependency graph. Otherwise, say ”Could not retrieve package type or version” and NO_ANSWER.
$\textrm{p}_{2}$ Phase-2 Task Instructions: After constructing the dependency graph successfully, you will have access to the Neo4j graph database, which contains the dependency graph. You will try your best to answer my questions. Note that:
1. You can use ”GraphSchemaTool” to get node labels and relationships in the dependency graph.
2. You can use ”CypherQueryTool” to get relevant information from the graph database. I will execute this query and send you the result. Make sure your queries comply with the database schema.
3. Use the ”WebSearchTool” to get information if needed.
4. To show the dependency graph, use ”VisualizeKGTool”.
Figure 2. Dependency KG Schema consisting of a single entity with two attributes, name and version, and one relation.
Finally, we equipped DepsRAG with utility tools, such as VisualizeKGTool, which is used to visualize the KG. Table 1 summarizes the tools used by DepsRAG.
Figure 3. Phase-1 of DepsRAG in action.
Figure 4. Phase-2 of DepsRAG in action.
Table 1. Summary of tools used by DepsRAG.
Tool Name Retriever Purpose
GraphSchemaTool KG Obtain the KG schema
CypherQueryTool KG
Receive the generated query, execute it,
and return the query result
WebSearchTool Web Search
Perform Web search, receive top-3 results,
and return the most relevant
ConstructKGTool N/A Construct the KG
VisualizeKGTool N/A Visualize the KG
4. Proof-of-Concept
This section describes the implementation and workflow of DepsRAG. We also discuss its capabilities, such as planning and multi-step tasks.
4.1. Implementation
We implemented DepsRAG in Python using Langroid (lan, 2024), which is a framework that supports the development of multi-agent LLMs and enables seamless integration with various LLMs. Moreover, Langroid can orchestrate the interaction between LLMs and tools, and contains built-in tools that facilitate the development of RAG applications, such as performing Web search and accessing Neo4j (neo, [n. d.]), a graph database with a declarative query language called Cypher.
To obtain the dependencies for the provided project and construct the dependency KG, we used Deps.Dev API, [^3] https://deps.dev a service developed and hosted by Google to help developers understand the structure and security of open-source software packages. This service was used in previous work (Rahman et al., 2024; Lin et al., 2024) to characterize and analyze dependencies. It repeatedly examines websites, such as GitHub and PyPI, to find up-to-date information about open-source software packages, thus generating a comprehensive list of direct and transitive dependencies that are constantly updated. DepsRAG receives a JSON response from the Deps.Dev API that contains a list of dependencies of the provided package. Then ConstructKGTool propagates the data in the JSON data to build the dependency KG. We have chosen two representative models to test DepsRAG, summarized in Table 2. Both are pre-trained LLMs. GPT-4 Turbo represents a proprietary LLM, whereas Llama-3 is an open-source model.
Table 2. Characteristics of the selected LLMs
Model Version Cut-off
Context Window
Input Cost
per 1M tokens
Output Cost
per 1M tokens
GPT-4 Turbo 1106-preview Apr 2023 128k \$10 \$30
Llama-3 70b-instruct Dec 2023 8k N/A N/A
Figure 5. Dependency graphs generated by DepsRAG for TensorFlow across 3 ecosystems (PyPI, NPM, and Rust). Versions 2.16.1, 0.7.0, and 0.21.0, respectively.
4.2. Capabilities Evaluation
  • Software Ecosystems Support. DepsRAG supports the construction of dependency graphs for four software ecosystems: PyPI, NPM, Cargo, and Go. Figure 5 illustrates the dependency graphs of the TensorFlow package in three ecosystems.
  • Cypher Query Generation. The performance of the KG Retriever depends on the model’s capability to translate a user query from the natural language to Cypher queries. We evaluated the two models used by DepsRAG to answer questions about important properties of graphs (i.e., depth, diameter, and transitivity). We found that both models generate correct Cypher queries, but not every time. To mitigate this situation, we instruct the LLMs to obtain the schema before generating the query and send the error message to the LLM to retry again. Table 3 (3rd column) indicates that Llama-3 failed to generate correct Cypher queries from the first trial for two questions. It made a single attempt to generate the query for the first question and had two attempts for the second question. While GPT-4-Turbo successfully produced accurate Cypher queries for all questions. Nonetheless, further analysis is necessary to assess this matter.
  • Task Planning and Decomposition. This is vital for breaking tasks into manageable sub-queries because LLMs often struggle with multi-step reasoning. To evaluate this aspect, we initially asked both models: “tell me two interesting things about the dependency graph.” Both demonstrated the capability to generate multiple Cypher queries, receive the answers and provide a complete answer. Table 3 provides examples for user questions and the corresponding Cypher queries, along with the responses generated by the considered model. The last question provides an additional example to demonstrate the capability of DepsRAG to perform task planning. The computation of graph density involves breaking this task into two steps: (1) retrieving the number of nodes (N) and edges (E) and (2) calculating the density according to this formula $\frac{E}{N(N-1)}$, since the KG is a directed graph. Llama-3 performed both steps but failed to give a correct answer as it assumed that the KG is an undirected graph and applied the incorrect formula.
Table 3. Analysis of dependency graph characteristics
Question # of Steps Query Trials Model Correct Response
What is the depth of the graph? 1 0 GPT-4-Turbo Yes
1 1 Llama-3 Yes
Are there cycles in the graph? 1 0 GPT-4-Turbo Yes
1 0 Llama-3 Yes
How many path chains are in the graph? 1 0 GPT-4-Turbo Yes
1 2 Llama-3 No
What is the density of the graph? 2 0 GPT-4-Turbo Yes
2 0 Llama-3 No
5. Related Work
Bommarito et al. (Bommarito and Bommarito, 2019) analyzed PyPI, including both package metadata and package source, covering software dependencies, authors, licenses, and other summary information over time. Musco et al. (Musco et al., 2017) constructed dependency graphs for Java programs at the class level to understand commonalities between different Java projects. In contrast, DepsRAG supports four software ecosystems and constructs dependency graphs at the package level.
Litzenberger et al. (Litzenberger et al., 2023) proposed a unified data model to implement and construct dependency graphs for arbitrary repositories, thus facilitating the comparison of dependency graphs between different repositories. In their implementation, the data model is based on Neo4j, which makes is compliant with DepsRAG’s KG schema. Maninger et al. (Maninger et al., 2023) proposed a visionary approach involving KGs to create a trustworthy AI software development assistant. Specifically, KGs can enable LLMs to correctly and appropriately explain the generated code. To our knowledge, we are the first to investigate the ability of LLMs to generate KGs for software dependencies, and utilize these graphs to aid in responding to queries concerning the dependency structure.
6. Conclusion and Future Work
We presented the first steps toward automatically generating KGs for software dependencies in four software ecosystems. There are several directions in which we plan to continue exploring this topic:
  • Extending Dependencies. The current KG is constructed based on Deps.Dev API that suffers from some limitations. For example, Python allows defining optional dependencies, indicating these dependencies will not be installed by default. However, Deps.Dev API does not capture these optional dependencies. This lack of support can lead to serious challenges and conflicts after updating packages that depend on optional packages. Therefore, retrieving dependencies from code and other dependency configuration files, such as “pyproject.toml,” is crucial. Moreover, certain software ecosystems may incorporate dependencies originating from different ecosystems. Previous studies have demonstrated the widespread utilization of C libraries within Python and NPM projects (Xu et al., 2022). Consequently, the dependency KG can be expanded to encompass these relationships, distinguishing between inter-ecosystem and intra-ecosystem dependencies.
  • Generating Software Bill of Materials (SBOM). The generated dependency graph encompasses both direct and transitive dependencies, enhancing the suitability of DepsRAG for SBOM generation and delivering metadata like path chains (i.e., dependency hierarchy (Balliu et al., 2023)), as opposed to presenting a flat list of dependencies. SBOM format specifications, such as CycloneDX (cyc, [n. d.]), are designed to accommodate this form of structured data. The hierarchical structure of the dependencies upheld by DepsRAG is in accordance with the minimum SBOM requirements prescribed by the US National Telecommunications and Information Administration (NTIA). DepsRAG exceeds these regulatory requirements, which require disclosure of primary dependencies (first-level) and all ensuing transitive dependencies (second-level) (nti, [n. d.]), by documenting the complete chain of dependencies. The generation of SBOM will enable DepsRAG to support vulnerability management, which requires designing a dedicated retrieval to access different sources of vulnerability databases.
  • Dependency Resolution. Updating dependencies is cumbersome and can introduce incompatibilities that break the application (Dann et al., 2023). DepsRAG maintains knowledge about all dependencies within the application, thus qualifying it to provide recommendations that suggest non-vulnerable package versions and will not lead to compatibility issues.